Calculate the limiting reagent in the following problem. When 25.0g of P4 reacts with 50.0g of O2, How much P4O10 is formed? P4 + 5 O2=>
P4O10
1
Expert's answer
2020-03-31T13:43:18-0400

"P_4 +5O_2\longrightarrow P_4O_{10}"


Here moles of "P_4=\frac{given mass}{molar mass}=\frac{25}{124}=0.201" mole


moles of "O_2= \frac{given mass}{molar mass}=\frac{50}{32}=1.5625" mole


here from given reaction 1 mole of P4 react with 5 mole of O2


so 1.5625 mole of O2 react with "=\frac{1}{5}\times1.5625=0.3125" mole of P4


So here limiting reagent is P4


Mole of "P_4O_{10}" produce from 0.201 mole of P4= 0.201 mole


So, mass of P4O10 produce= "mass\times molar mass" = "0.201\times" 283.88=56.94 g


Mass of P4O10 produce= 56.94 g